Solve dy/dx-4y=2x^2 using the method exploit the superstition.


1
Expert's answer
2022-02-01T10:53:34-0500
"y=uv"

"y'=u'v+uv'"

Substitute


"u'v+uv'-4uv=2x^2"

"uv'+v(u'-4u)=2x^2"

Put the "v"  term equal to zero


"u'-4u=0"

"\dfrac{du}{u}=4dx"

Integrate


"\int \dfrac{du}{u}=\int4dx"


"u=Ce^{4x}"

Then


"Ce^{4x}v'=2x^2"

"dv=\dfrac{2}{C}x^2 e^{-4x}dx"

Integrate


"\int dv=\int\dfrac{2}{C}x^2 e^{-4x}dx"

"v=-\dfrac{1}{2C}x^2e^{-4x}-\dfrac{1}{4C}xe^{-4x}-\dfrac{1}{16C}e^{-4x}+C_1"

"y=-\dfrac{1}{2}x^2-\dfrac{1}{4}x-\dfrac{1}{16}+C_2e^{4x}"